# Large language model validity via enhanced conformal prediction methods

### v1

@@ 3.3 Conditional boosting
In this section, we describe a new method for automated conformity score design subject to conditional coverage constraints. As discussed in the introduction, the goal of this procedure is to find conformity scores that when combined with our filtering method, allow the filter to retain as much of the LLM output as possible while still ensuring validity. To make our exposition more concrete, we will assume for the sake of simplicity that we are boosting the conformity score defined in (6). Our approach, however, will generalize to any parameterized score function.
@@
There are therefore two obstacles to optimizing (9) via gradient descent. First, the indicator function is non-differentiable. To resolve this, we approximate the indicator using a sigmoid function [25]. Second, it is unclear how to backpropagate through the estimated cutoff $\hat{\tau}_{i}(\theta)$.
Observe that for linear function classes =$\mathcal{F}=\{\Phi(X)^{\top}\beta:\beta\in\mathbb{R}^{d}\}$, the regression (8) by which we obtain $\hat{\tau}_{i}(\theta)$ is a linear program. Using this observation, we find that $\hat{\tau}_{i}(\theta)$ can be expressed as the solution to a linear system of equations given by a subset of the dataset that we denote by $B$, i.e.,
$\hat{\tau}_{i}(\theta)=\Phi(X)_{B}^{-1}S_{B}(\theta).$ (10)
In the linear programming literature, this subset is typically referred to as the “optimal basis.” It can be explicitly identified by selecting the points at which the quantile regression interpolates the scores. Then, so long as $B$ is invariant to small perturbations of $\theta$ and $S$ is differentiable in $\theta$, we can obtain the derivative of $\hat{\tau}_{i}(\theta)$ with respect to $\theta$ by backpropagating through (10). Proposition 3.1 identifies a simple condition under which this is possible.
@@
Assume that the solution to the quantile regression given by (7) with $S=\hat{\tau}_{i}(\theta)$ is non-degenerate, i.e., it interpolates exactly $\text{dim}(\Phi)$ points, and let $B$ denote the optimal basis of this solution. Then, $\hat{\tau}_{i}(\theta)$ has derivative
$\partial_{\theta}\hat{\tau}_{i}(\theta)=\Phi(X)_{B}^{-1}\partial_{\theta}S_{B}(\theta).$ (11)
The non-degeneracy assumption in Proposition 3.1 nearly always holds; if it does not, we remark that non-degeneracy can always be obtained by “dithering” the scores, i.e., adding small i.i.d. noise, at each step of the boosting procedure [7].
Now that we have resolved the gradient computation problem, the remainder of our procedure emulates the analogous ConfTr algorithm of Stutz et al. [25]. Similar to the $\alpha(\cdot)$-estimation problem, we must leave a portion of the dataset for the final calibration procedure. In our experiments, we re-use the portion set aside for boosting. With this part of the available data, we re-split our dataset at each iteration into a calibration and test set, compute the quality criterion on the test set, and backpropagate on the objective given by (9) subject to the above modifications. For the sake of space, we defer Algorithm 1, which provides a complete description of the procedure, to Appendix C.
@@ 4.1 Synthetic data
Figure 3: Comparison of the marginal boosting procedure of Stutz et al. [25] (blue) against our conditional boosting method (orange). The left panel shows the prediction sets produced by each method, while the right panel displays the conditional coverage, $\mathbb{P}(Y_{n+1}\in\hat{C}(X_{n+1})\mid X_{n+1})$ against the values of the first feature, $X_{n+1}^{(1)}$. Using the Adam optimizer with learning rate set to $0.001$, the plotted scores are boosted for $500$ steps on a synthetic dataset of size $n=1000$ and evaluated on another test dataset of size $n=2000$.
@@
Figure 3 compares the ConfTr algorithm of Stutz et al. [25] against our Conditional Boosting method. While the left panel shows that the former method issues shorter intervals, these come at the cost of degraded conditional validity (right panel). In particular, the prediction sets resulting from ConfTr have extremely poor coverage on the subset of data with high conditional variance. By contrast, learning $\theta$ using our procedure still attains near-perfect conditional coverage.
@@ 4.2 Medical long-form question-answering
In this section, we summarize the experimental setup for the claims and figures presented in Section 1.1. Our experiment considers the long-form medical question-answering dataset (MedLFQA) published in Jeong et al. [13]. It combines several previously established benchmarks in the medical question-answering literature: HealthSearchQA ($n=3047$), K-QA ($n=1077$), LiveQA ($n=100$) and MedicationQA ($n=627$). Each prompt in these datasets is also accompanied by either an LLM or human-generated response to each question [6, 1, 17, 24]. Following prior work [17, 13], we treat these responses as ground-truth for model evaluation. Also matching Jeong et al. [13], in our plots, we distinguish between the subset of questions ($n=201$) in the K-QA dataset that have gold-standard physician responses (denoted by K-QA Golden) and those questions with LLM-generated responses (denoted by K-QA Silver). To reproduce our results, we point the interested reader to our GitHub repository, which includes a filtered and combined dataset that removes some non-health-related prompts contaminating the HealthSearchQA dataset as well as some extraneous text from the questions.
To obtain the dataset used in our experiment, we query GPT-3.5-Turbo with each of the prompts in the combined dataset, and then ask GPT-4o to parse these prompts into self-contained sub-claims. Then, to obtain our “ground-truth” labels, we ask GPT-3.5-Turbo to evaluate whether each claim is substantiated by the pseudo-ground-truth response for that prompt. The prompts we use for each of these tasks are included in Section D.2.
@@
On the first split ($n=1421$), we boost our claim score via optimizing a linear combination of four scores previously described in the literature: the frequency, self-evaluation, and ordinal scores described in Mohri and Hashimoto [19] and the token-level probability assigned to a single-token self-evaluation [14]. Section D.1 includes a precise description of how we compute these features. We then differentiate through the conformal calibration step with respect to the function class $\mathcal{F}$ defined by the linear combination of an intercept, the number of characters in the prompt, the number of characters in the response, the mean frequency score assigned to the claims, and the standard deviation of the frequency scores assigned to the claims. The results shown in Figure 2 are obtained by running the conditional boosting algorithm for $1000$ steps using the Adam optimizer with learning rate set to $0.001$.
Having obtained an improved conformity score, we simulate the conditional conformal step and estimate $\alpha(\cdot)$ by dividing this portion of the dataset in half. On the first half of this dataset ($n=720$), we run the fixed-level conditional conformal procedure (with the same $\mathcal{F}$ used to boost the scores) for each $\alpha$ in a grid of $50$ evenly spaced values between 00 and $1$. Using the second half of this dataset, we evaluate the percentage of retained claims at each $\alpha$. We record the smallest $\alpha$ for which at least $70\%$ of claims are preserved at all larger values of $\alpha$. We then obtain our estimate of $\alpha(\cdot)$ by fitting a $0.15$-quantile regression over $\mathcal{F}$ to this dataset. For interpretability and numerical stability in the next step, the final output of the function is truncated to lie between $0.1$ and $0.5$.
On the second half of this dataset, we run level-adaptive conformal prediction over the previously described $\mathcal{F}$ augmented by indicator functions that correspond to $\alpha(\cdot)$ falling in some sub-interval with endpoints lying in $\{0,0.05,\dots,1\}$. Finally, to produce the plots in Figure 2, we run this step over $100$ random calibration-test splits of size $2354$ and $500$, respectively.
@@ 4.3 Wikipedia biographies
Figure 5: Empirical demonstration of our method on the Wikipedia biographies dataset. The left two panels display results for our level-adaptive method, which aims to issue biographies with 3 or fewer errors, while retaining at least 80% of the original claims in the prompt. The left panel compares the binned nominal probabilities reported by our method with bin width $2.5\%$ against the true realized empirical values evaluated over $400$ test points and 100 calibration-test splits. The center panel compares the number of claims retained by our method (orange) against the fixed level method of Mohri and Hashimoto [19] (blue). In this plot, the $y$-axis displays a moving average of the number of claims retained with window size $1000$, while the $x$-axis a moving average of the number of views (in Jan. 2023) of the Wikipedia article associated with each prompt on the log-scale. Finally, the right-hand panel displays the claim retention obtained with boosted (orange) and unboosted (blue) scores at a fixed level of $\alpha=0.1$. Boxplots in this panel show the distribution of retentions for 100 calibration-test splits each containing 7246 calibration points and 381 test points.
Our LLM experiment reconsiders the biography dataset analyzed in Mohri and Hashimoto [19]. We sample $8516$ names from Wikipedia and query GPT-3.5-Turbo with the prompt “Write me a short biography of [NAME].” We then ask GPT-4o to parse the prompts into self-contained sub-claims. While the prior work only annotates $25$ biographies, we need to scale up this experiment to validate our proposed methods. To get around expensive human annotation, we use a variant of the FActscore procedure developed by Min et al. [18]. Their method, commonly known as retrieval-augmented generation (RAG), passes in relevant Wikipedia passages identified by the BM25 ranking function as part of the prompt to the LLM and asks if the claims are supported. Scoring of the LLM outputs (i.e., computation of $p(P_{i},C_{ij})$) is done using several self-evaluation methods previously described in the LLM alignment literature [19, 14]. A more detailed description of the methods described in this paragraph can be found in Sections D.1 and D.2.
@@
We control this relaxed factuality criterion so that we can issue non-trivial guarantees over all topics. For the sake of avoiding redundancy with the previous subsection, further details about how our methods are instantiated are deferred to Section D.3.
Figure 6: Comparison of the split conformal calibration method of Mohri and Hashimoto [19] (blue) against our conditional conformal method (orange). The left and right panels displays the miscoverage and percentage of claims retained by the two methods against the number of views received by the Wikipedia pages in January 2023. The displayed boxplots are computed over $200$ trials in which we run both methods on a calibration set of $5890$ points and evaluate their coverage on a test set of size $2500$. The displayed groups correspond to view counts that are binned into the intervals (,)($\infty$,$\infty$)( - ∞ , ∞ ), $[1000000,\infty)$, $[100000,1000000)$, $[1000,100000)$, $[100,1000)$, $[0,100)$
@@ Appendix A Additional details about conditional conformal
We note that Gibbs et al. [10] also consider extensions of Theorem 2.1 to other function classes and regularized quantile regression procedures. All of these extensions can be readily applied to our present setting.
Formally, Gibbs et al. [10] consider extensions of their procedure to cases where $\mathcal{F}$ is a general vector space and $g_{S}$ is fit using a penalized regression. In this section, we extend the work of Gibbs et al. [10] to our current setting. We then describe the randomization procedure that is used to construct our filtering threshold, $\hat{\tau}_{\text{l.a. rand.}}(X_{n+1})$ and prove Theorems 3.1 and 3.2.
To extend our procedure to general function classes and regularizers, we consider regressions of the form
$g_{S}=\underset{g\in\mathcal{F}}{\text{argmin}}\frac{1}{n+1}\sum_{i=1}^{n}\ell_{\alpha(X_{i})}(S(\mathbf{C}_{i},\mathbf{W}_{i})-g(X_{i}))+\frac{1}{n+1}\ell_{\alpha(X_{n+1})}(S-g(X_{n+1}))+\mathcal{P}(g),$ (12)
where $\mathcal{P}(\cdot)$ denotes any convex penalty. Following our work in the main text, this leads to the filtering threshold $\hat{\tau}_{\text{l.a.}}(X_{n+1})=\max\{S:S\leq g_{S}(X_{n+1})\}$. As discussed earlier, this definition of $\hat{\tau}_{\text{l.a.}}(X_{n+1})$ can be slightly conservative and thus it can be preferrable to work with a randomized analog. We derive this analog now. First, note that (12) is equivalent to the program
$\begin{split}\underset{p,q\in\mathbb{R}^{n+1},\ g\in\mathcal{F}}{\text{minimize}}\quad&\sum_{i=1}^{n+1}(1-\alpha)p_{i}+\alpha q_{i}+(n+1)\mathcal{P}(g),\\ \text{subject to}\quad&S(\mathbf{C}_{i},\mathbf{W}_{i})-g(X_{i})-p_{i}+q_{i}=0,\ \forall 1\leq i\leq n,\\ &S-g(X_{n+1})-p_{n+1}+q_{n+1}=0,\\ &p_{i},q_{i}\geq 0,\ \forall 1\leq i\leq n+1.\end{split}$
@@
Proof   [Proof of Theorem A.1] By the monotonicty assumption on $L(\cdot,\cdot)$ and our definition of the conformity score, we have that
@@
Proof   [Proof of Theorem 3.2] This result is the special case of Theorem A.1 in which $\mathcal{P}(\cdot)=0$, $\alpha(X_{i})=\alpha$ is a constant function, and $\mathcal{F}$ is a finite-dimensional linear class. Under this setup, the primal and dual programs outlined above are linear program. Since the dual program admits the interior feasible solution $\eta=0$, Slater’s condition immediately implies that these programs satisfy strong duality. ∎

### v2

@@ 3.3 Conditional boosting
In this section, we describe a new method for automated conformity score design subject to conditional coverage constraints. As discussed in the introduction, the goal of this procedure is to find conformity scores that when combined with our filtering method, allow the filter to retain as much of the LLM output as possible while still ensuring validity. For the sake of simplicity, we will assume that we are boosting the conformity score defined in (6). Our approach, however, will generalize to any parameterized score function. We note here that the concurrent work of Kiyani et al. [16] presents another approach to this problem by reframing it as a min-max optimization task.
@@
There are therefore two obstacles to optimizing (9) via gradient descent. First, the indicator function is non-differentiable. To resolve this, we proceed as in Stutz et al. [27] and approximate the indicator using a sigmoid function. Second, it is unclear how to backpropagate through $\hat{\tau}_{i}(\theta)$.
Observe that for a linear function class =$\mathcal{F}=\{\Phi(X)^{\top}\beta:\beta\in\mathbb{R}^{d}\}$, the regression (8) by which we obtain $\hat{\tau}_{i}(\theta)$ is a linear program. Using this observation, we find that $\hat{\tau}_{i}(\theta)$ can be expressed as the solution to a linear system of equations given by a subset of the dataset that we denote by $B$, i.e.,
$\hat{\tau}_{i}(\theta)=\Phi(X_{n+i})^{\top}\left(\Phi(X)_{B}^{-1}S_{B}(\theta)\right).$ (10)
Here, $\Phi(X)_{B}^{-1}$ and $S_{B}(\theta)$ are used to denote the matrix of features and vector of scores for the subset of the data given by $B$. In the linear programming literature, this subset is typically referred to as the “optimal basis.” It can be explicitly identified by selecting the points at which the quantile regression interpolates the scores. Then, so long as $B$ is invariant to small perturbations of $\theta$ and $S$ is differentiable in $\theta$, we can obtain the derivative of $\hat{\tau}_{i}(\theta)$ with respect to $\theta$ by backpropagating through (10). Proposition 3.1 identifies a simple condition under which this is possible.
@@
Let $\hat{\tau}_{i}(\theta)$ denote the non-randomized filtering threshold defined via (8). Assume that the threshold is finite and that the augmented quantile regression for all $S>\hat{\tau}_{i}(\theta)$ admits a unique non-degenerate basic solution. Let $B$ denote this basis. Then, $\hat{\tau}_{i}(\theta)$ has derivative
$\partial_{\theta}\hat{\tau}_{i}(\theta)=\Phi(X_{n+i})^{\top}\left(\Phi(X)_{B}^{-1}\partial_{\theta}S_{B}(\theta)\right).$ (11)
The technical condition on the basis in Proposition 3.1 nearly always holds; if it does not, we remark that uniqueness can always be obtained by “dithering” the scores and/or design matrix, i.e., adding small i.i.d. noise, at each step of the boosting procedure [7].
The rest of our procedure emulates the ConfTr algorithm of Stutz et al. [27]. Having set aside a portion of the dataset for the final calibration procedure, we replicate the conformal prediction algorithm at each iteration. Namely, we randomly split the remaining data into a “calibration” and “test” set. We then run the level-adaptive conformal prediction method on the calibration set, compute the quality criterion on the test set, and backpropagate on the objective given by (9). Algorithm 1, which presents a complete description of the procedure, can be found in Appendix C.
@@ 4.1 Synthetic data
Figure 3: Comparison of the marginal boosting procedure of Stutz et al. [27] (blue) against our conditional boosting method (orange). The left panel shows the prediction sets produced by each method, while the right panel displays the conditional coverage, $\mathbb{P}(Y_{n+1}\in\hat{C}(X_{n+1})\mid X^{(1)}_{n+1})$ against the values of the first feature, $X_{n+1}^{(1)}$. Using the Adam optimizer with learning rate set to $0.001$, the plotted scores are boosted for $500$ steps on a synthetic dataset of size $n=1000$ and evaluated on another test dataset of size $n=2000$.
@@
Figure 3 compares the ConfTr algorithm of Stutz et al. [27] against our Conditional Boosting method. While the left panel shows that the former method may sometimes issue shorter intervals, these come at the cost of degraded conditional validity (right panel). In particular, the prediction sets resulting from ConfTr have extremely poor coverage on the subset of data with high conditional variance. By contrast, learning $\theta$ using our procedure still attains near-perfect conditional coverage.
@@ 4.2 Medical long-form question-answering
In this section, we summarize the experimental setup for the claims and figures presented in Section 1.1. Our experiment considers the long-form medical question-answering dataset (MedLFQA) published in Jeong et al. [13]. It combines several previously established benchmarks in the medical question-answering literature: HealthSearchQA ($n=3047$), K-QA ($n=1077$), LiveQA ($n=100$) and MedicationQA ($n=627$). Each prompt in these datasets is also accompanied by either an LLM or human-generated response to each question [6, 1, 19, 26]. Following prior work [19, 13], we treat these responses as ground-truth for model evaluation. Also matching Jeong et al. [13], in our plots, we distinguish between the subset of questions ($n=201$) in the K-QA dataset that have gold-standard physician responses (denoted by K-QA Golden) and those questions with LLM-generated responses (denoted by K-QA Silver). To reproduce our results, our GitHub repository includes a filtered and combined dataset that removes some non-health-related prompts contaminating the HealthSearchQA dataset.
To obtain the dataset used in our experiment, we query GPT-3.5-Turbo with each of the prompts in the combined dataset, and then ask GPT-4o to parse these prompts into self-contained sub-claims. Then, to obtain our “ground-truth” labels, we ask GPT-3.5-Turbo to evaluate whether each claim is substantiated by the pseudo-ground-truth response for that prompt. The prompts we use for each of these tasks are included in Section E.2.
@@
On the first split ($n=1421$), we boost our claim score and subsequently estimate the level required for $70\%$ claim retention. We achieve the former by optimizing a linear combination of four scores previously described in the LLM factuality literature: the frequency, self-evaluation, and ordinal scores described in Mohri and Hashimoto [21] and the token-level probability assigned to a single-token self-evaluation [14]. Section E.1 describes how these claim scores are computed. The function class $\mathcal{F}$ for which we run the conditional boosting method is defined by the linear combination of several prompt-response features. For the sake of brevity, we defer a detailed description of this function class, our method for estimating $\alpha(\cdot)$, and ablations of our method to Section D.1.
On the second half of this dataset, we run level-adaptive conformal prediction over the $\mathcal{F}$ used in the conditional boosting step augmented by indicator functions that correspond to $\alpha(\cdot)$ falling in some sub-interval with endpoints lying in $\{0,0.05,\dots,1\}$. Finally, to produce the plots in Figure 2, we run this step over $100$ random calibration-test splits of size $2354$ and $500$, respectively.
@@ 4.3 Wikipedia biographies
Figure 5: Empirical demonstration of our method on the Wikipedia biographies dataset. The left two panels display results for our level-adaptive method, which aims to issue biographies with 3 or fewer errors, while retaining at least 80% of the original claims in the prompt. The left panel compares the binned nominal probabilities reported by our method with bin width $2.5\%$ against the true realized empirical values evaluated over $381$ test points and 100 calibration-test splits. The center panel compares the number of claims retained by our method (orange) against the fixed level method of Mohri and Hashimoto [21] (blue). In this plot, the $y$-axis displays a moving average of the number of claims retained with window size $1000$, while the $x$-axis a moving average of the number of views (in Jan. 2023) of the Wikipedia article associated with each prompt on the log-scale. Finally, the right-hand panel displays the claim retention obtained with boosted (orange) and unboosted (blue) scores at a fixed level of $\alpha=0.1$. Boxplots in this panel show the distribution of retentions for 100 calibration-test splits each containing 7246 calibration points and 381 test points.
Figure 6: Comparison of the split conformal calibration method of Mohri and Hashimoto [21] (blue) against our conditional conformal method (orange). The left and right panels displays the miscoverage and percentage of claims retained by the two methods against the number of views received by the Wikipedia pages in January 2023. The displayed boxplots are computed over $200$ trials in which we run both methods on a calibration set of $5890$ points and evaluate their coverage on a test set of size $2500$. The displayed groups correspond to view counts that are binned into the intervals (,)($\infty$,$\infty$)( - ∞ , ∞ ), $[1000000,\infty)$, $[100000,1000000)$, $[1000,100000)$, $[100,1000)$, $[0,100)$. The fraction of the data set belonging to each group (in plotted order) is $8.7\%,30.9\%,39.5\%,19.3\%,$ and $1.5\%$, respectively.
Our second experiment reconsiders the biography dataset analyzed in Mohri and Hashimoto [21]. We sample $8516$ names from Wikipedia and query GPT-3.5-Turbo with the prompt “Write me a short biography of [NAME].” We then ask GPT-4o to parse the prompts into self-contained sub-claims. While the prior work only annotates $25$ biographies, we need to scale up this experiment to validate our proposed methods. To get around expensive human annotation, we use a variant of the FActscore procedure developed by Min et al. [20]. Their method includes relevant Wikipedia passages identified by the BM25 ranking function in the prompt to the LLM and asks if the claims are supported. Scoring of the LLM outputs (i.e., computation of $p(P_{i},C_{ij})$) is done using the frequency scoring method in all displayed figures except for the boosting comparison (the right panel of Figure 5). In that plot, we compare the claim retention achieved by an optimal ensemble of three cheaper-to-compute scores (self-evaluation, ordinal, log-probability) to the retention of a uniform mixture of those scores [21, 14]. A detailed description of the claim scoring and data collection methods mentioned in this paragraph can be found in Sections E.1 and E.2. Method ablations and experiments comparing the efficacy of each claim scoring approach are included in Section D.2.
@@
Empirically, we find that the claim scoring methods we apply are not as well-correlated with factuality as they were for MedLFQA. As a result, in order to issue non-trivial guarantees, i.e., to ensure both that $1-\alpha(X_{n+1})$ remains reasonably large and that the method does not filter too much of the response, we allow the filter to keep up to three false claims. [^4] While a non-zero error guarantee is not always desirable, it may be suitable for tasks where the cost of an error is low and/or high error rates appear unavoidable, e.g., “Please suggest ten recommendations for my vacation.” or “Give me a list of already-approved drugs that might be repurposed to treat COVID.” Experimental results for the more typical 00-error guarantee are also included in Section D.2. For the sake of avoiding redundancy with the previous subsection, we defer the remaining experimental details to Section D.2.
@@
5 Limitations
While we believe the described methodology can improve LLM factuality, we highlight some limitations here. First, our theoretical results assume that the prompt-response tuples are i.i.d. In settings where user interactions change over time, the test prompts may be incomparable to previous prompts. Even though our framework can be applied to guarantee validity under a pre-specified class of covariate shifts (see Section A.2 for details), robustness guarantees under other types of distribution shift will require additional research. Second, since our filter is a wrapper around existing claim scores and thus the utility of our method depends on the quality of the underlying scoring algorithm. If the claim scores are weakly correlated with factuality, the filtered output will be accompanied by a vacuous nominal guarantee. Nevertheless, discovering new features that accurately predict LLM hallucinations is an active field of research, and our wrapper can easily leverage new approaches.
Acknowledgments
E.J.C. was supported by the Office of Naval Research grant N00014-24-1-2305, the National Science Foundation grant DMS-2032014, and the Simons Foundation under award 814641. I.G. was also supported by the Office of Naval Research grant N00014-24-1-2305 and the Simons Foundation award 814641, as well as additionally by the Overdeck Fellowship Fund. J.J.C. was supported by the John and Fannie Hertz Foundation. The authors are grateful to Anav Sood and Tim Morrison for helpful discussion on this work.
@@ Appendix A Additional details about conditional conformal
In addition to the methods discussed in Section 2.1, Gibbs et al. [10] also consider extensions of Theorem 2.1 to other function classes and regularized quantile regression procedures. In this section, we show how all of these results can be applied to our present setting. We then describe the randomization procedure that is used to construct our filtering threshold, $\hat{\tau}_{\text{l.a. rand.}}(X_{n+1})$ and prove Theorems 3.1 and 3.2.
Following Gibbs et al. [10], let $\mathcal{F}$ be a vector space and $\mathcal{P}:\mathcal{F}\to\mathbb{R}$ be a convex penalty. Then, to extend our procedure to such function classes and regularizers, we consider regressions of the form
$g_{S}=\underset{g\in\mathcal{F}}{\text{argmin}}\frac{1}{n+1}\sum_{i=1}^{n}\ell_{\alpha(X_{i})}(S(\mathbf{C}_{i},\mathbf{W}_{i})-g(X_{i}))+\frac{1}{n+1}\ell_{\alpha(X_{n+1})}(S-g(X_{n+1}))+\mathcal{P}(g).$ (12)
Following our work in the main text, this leads to the filtering threshold $\hat{\tau}_{\text{l.a.}}(X_{n+1})=\max\{S:S\leq g_{S}(X_{n+1})\}$. As discussed earlier, this definition of $\hat{\tau}_{\text{l.a.}}(X_{n+1})$ can be slightly conservative and thus it can be preferrable to work with a randomized analog. We derive this analog now. First, note that (12) is equivalent to the program
$\begin{split}\underset{p,q\in\mathbb{R}^{n+1},\ g\in\mathcal{F}}{\text{minimize}}\quad&\sum_{i=1}^{n+1}(1-\alpha(X_{i}))p_{i}+\alpha(X_{i})q_{i}+(n+1)\mathcal{P}(g),\\ \text{subject to}\quad&S(\mathbf{C}_{i},\mathbf{W}_{i})-g(X_{i})-p_{i}+q_{i}=0,\ \forall 1\leq i\leq n,\\ &S-g(X_{n+1})-p_{n+1}+q_{n+1}=0,\\ &p_{i},q_{i}\geq 0,\ \forall 1\leq i\leq n+1.\end{split}$
@@
Proof   [Proof of Theorem A.1] By Theorem 4 of Gibbs et al. [10], $S\mapsto\eta^{S}_{n+1}$ is non-decreasing in $S$. So, by the monotonicity assumption on $L(\cdot,\cdot)$ and our definition of the conformity score, we have that
@@
Proof   [Proof of Theorem 3.1] This result is the special case of Theorem A.1 in which $\mathcal{P}(\cdot)=0$, $\alpha(X_{i})=\alpha$ is a constant function, and $\mathcal{F}$ is a finite-dimensional linear class. Under this setup, the primal and dual programs outlined above are linear programs. Since the dual program admits the interior feasible solution $\eta=0$, Slater’s condition immediately implies that these programs satisfy strong duality. ∎
@@
A.2 Interpretation of the guarantees in terms of covariate shifts
As discussed in the original work of Gibbs et al. [10], the guarantees on the loss we developed in Theorems 3.1 and 3.2 can be readily interpreted as a form of robustness with respect to a class of distribution shifts affecting $P_{X}$ (but not $P_{Y\mid X}$) known as covariate shifts.
First, to precisely define such a shift, suppose that $f\in\mathcal{F}$ is non-negative. Recall that our features $X_{n+1}=X(P_{n+1},R_{n+1})$ are functions of the prompt and response. Consider the setting in which the training data, $\{(P_{i},R_{i},\mathbf{C}_{i},\mathbf{W}_{i})\}_{i=1}^{n}$ are sampled i.i.d. from a distribution $Q$, while the test point is sampled from the “covariate-shifted” distribution
$(P_{n+1},R_{n+1})\sim\frac{f(X_{n+1})}{\mathbb{E}_{Q}[f(X)]}dQ_{(P,R)},\qquad(\mathbf{C}_{n+1},\mathbf{W}_{n+1})\sim Q_{(\mathbf{C},\mathbf{W})|(P,R)}.$
Here, we use the notation $\frac{f(X_{n+1})}{\mathbb{E}_{Q}[f(X)]}dQ_{(P,R)}$ to refer to the distribution in which the covariate space is re-weighted by $f$, e.g. using the subscript $f$ to denote the re-weighting, we have that for any set $A$,
$\mathbb{P}_{f}(X_{n+1}\in A)=\frac{\mathbb{E}_{Q}[f(X)\mathbf{1}\{X\in A\}]}{\mathbb{E}_{Q}[f(X)]}.$
Then, using these definitions, our Theorems 3.1 and 3.2 have the following immediate corollaries.
Corollary A.1 (Extension of Theorem 3.1).
Let $\mathbb{P}_{f}$ denote the covariate shift setting described above. Then, under the setting and assumptions of Theorem 3.1,
$\mathbb{P}_{f}(L(\hat{F}(\mathbf{C}_{n+1};\hat{\tau}_{\textup{rand}}(X_{n+1})),\mathbf{W}_{n+1})\leq\lambda)=1-\alpha,\ \forall f\in\{h\in\mathcal{F}:0<\mathbb{E}_{Q}[h(X)]<\infty\}.$
Corollary A.2 (Extension of Theorem 3.2).
Let $\mathbb{E}_{f}[\cdot]$ denote expectations with respect to the covariate shift setting described above. Then, under the setting and assumptions of Theorem 3.2,
$\mathbb{E}_{f}[(\mathbbm{1}\{L(\hat{F}(\mathbf{C}_{n+1};\hat{\tau}_{\textup{l.a., rand.}}(X_{n+1})),\mathbf{W}_{n+1})\leq\lambda\}-(1-\alpha(X_{n+1})))]=0,$
$\hskip 56.9055pt\forall f\in\{h\in\mathcal{F}:0<\mathbb{E}_{Q}[h(X)]<\infty\}.$